OpenStudy (anonymous):

express using positive exponents then if possible simplify 1 over z^6

OpenStudy (anonymous):

\[\bf \frac{ 1 }{ x^n }=x^{-n}\]

OpenStudy (anonymous):

i cannot really tell what the number after x is

OpenStudy (anonymous):

\[\bf \frac{ 1 }{ z^6 } =..?\]Can you tell what 'x' is and what 'n' is? @snoop313

OpenStudy (anonymous):

i am not sure what steps to take to get my answer

OpenStudy (goformit100):

ok first show your steps.

OpenStudy (anonymous):

I am not sure how to get started on this what is my first step?